# Manifolds, Random Matrices and Spectral Gaps: The geometric phases of generative diffusion

### v4

@@ 1 Introduction
Generative diffusion models have revolutionized the fields of computer vision and generative modeling, achieving state-of-the-art performance on image generation (Ho et al.,, 2020; Song and Ermon,, 2019; Song et al.,, 2021) and video generation (Ho et al.,, 2022; Singer et al.,, 2022; Blattmann et al.,, 2023; Brooks et al.,, 2024). Generative diffusion models synthesize images through a stochastic dynamical denoising process. Experimental and theoretical arguments suggest that different features such as frequency modes and class labels are generated at different times during the process. For example, it has been shown that separation between isolated classes, as in the case of mixture of Gaussian models, happens at critical phase transition points of spontaneous symmetry breaking (speciation events) (Biroli et al.,, 2024). It is also well known that subspaces corresponding to different frequency modes emerge at different times of diffusion (Kingma and Gao,, 2024). This idea has been recently refined by (Kadkhodaie et al.,, 2024), who showed that diffusion models give rise to a local decomposition of the image manifold into a basis of geometry-adaptive harmonic basis functions. These decomposition phenomena cannot be directly explained in terms of critical phase transitions as they are fundamentally linear processes. In this paper, we will provide a precise theoretical analysis of the separation of subspaces for data defined on low dimensional linear manifolds.
@@ 3 Background on generative diffusion models
where $\text{d}\boldsymbol{Z}_{t}$ is a standard Brownian motion. The formal solution of Eq. (3) can be given in terms of the heat kernel $p(\boldsymbol{x}_{t},t)=\mathbb{E}_{\boldsymbol{x}_{0}\sim p_{0}}\!\left[\frac{1}{\sqrt{2\pi t}}e^{-\frac{\left\lVert\boldsymbol{x}_{t}-\boldsymbol{x}_{0}\right\rVert_{2}^{2}}{2t}}\right].$ The target distribution $p_{0}(\boldsymbol{x})$ is then recovered by reversing the diffusion process (Anderson,, 1982). We initialize this reverse process from $\boldsymbol{x}_{t_{f}}\sim\mathcal{N}\left(0,t_{f}I_{d}\right)$ at some large time $t_{f}$, which evolves backward in time according to
@@ 5.4 The geometric phases and manifold overfitting
The probability density of data defined on a manifold is a spiked object $\delta_{\mathcal{M}}(x)\rho(x)$, where the Dirac-delta $\delta_{\mathcal{M}}(x)$ determines the manifold and $\rho(x)$ determines its internal density. Likelihood-based generative models are defined by a highly parameterized likelihood function $f(\boldsymbol{x};\boldsymbol{\theta})$, whose parameters are trained by minimizing the loss
@@
which maximizes the probability of the data given the model. This maximum likelihood loss is minimized if $f(\boldsymbol{x};\boldsymbol{\theta})=p_{0}(\boldsymbol{x})$. A trained likelihood-based model can only fit the true density by having it to diverge to infinity on the manifold. Such a divergence makes it impossible to correctly model the internal density $\rho(x)$. More problematically, the optimization problem becomes almost insensitive to the internal density $\rho(\boldsymbol{x})$. This phenomenon is called manifold overfitting (Loaiza-Ganem et al.,, 2022), since the trained model fits the manifold while ignoring its internal density, resulting in poor generation.
Our analysis suggests that the temporal dynamics of generative diffusion models overcome this limitation because, for intermediate values of $t$, the score is still sensitive to the density internal to the manifold, which can be identified through the differences in the tangent singular values. During this manifold coverage phase, the score directs the dispersion of the particles according to these differences, with higher singular values resulting in larger ‘opposing force’ from the score, which results in smaller displacements of the generated samples along these directions. For $t$ tending to zero, these differences are suppressed due to the divergence of the likelihood, which results in a score function that is orthogonal to the manifold and that is insensitive to $\rho(x)$. However, at this stage of generative diffusion the internal dispersion of the particles have already been affected by the previous coverage phase and therefore the manifold overfitting of the score does not negatively affect generation. Instead, the consolidation phase plays the important role of projecting the particles to the support of the data.
@@ 6.1 Linear manifolds
6.2 The linear random-manifold score in large dimensions: A random matrix analysis
In the following, we provide an outline of our theoretical results on the distribution of spectral gaps in the matrix $J_{t}$ under random linear manifolds. This choice reflects the fact that the distribution and support of the data are usually not known in advance, and it is therefore important to quantify the statistical variability induced by this uncertainty. Specifically, we will consider random projection matrices $F$ distributed according to different distributions. To ensure tractability, we perform the analysis in the limit of large $d$ (visible) and $m$ (latent) dimensions while keeping the ratio $\alpha_{m}=m/d$ constant.
@@ 6.4 Intermediate gaps and subspaces with different variances
This same analysis can be extended to the more general case where the spectral density is known to be formed by different detached bulks, associated with hierarchically smaller variances of the data. The evolution of the intermediate gaps in a double-variance diffusion model is reported in Fig. 2: notice that $t_{\text{max}}=\mathcal{O}(\sigma_{2})$ is consistent with Fig. 2b and 2f, where the gap was found to be maximum in width. It is worth noting that subspaces with higher variances are the first ones to be explored by diffusion, and to be learned by the model. This point suggests that the model is sensitive to the parameters of the probability distribution on the manifold as recently suggested by other works in the literature (see Section 9 for further details).
@@ 8 Experiments with natural image datasets
  1. 1.
    Trivial phase: at large times (i.e. from $t=200$ to $t=100$) the ordered spectrum of the singular values appears flat, suggesting the diffusive motion to be Brownian in the ambient space.
  2. 2.
@@ 9 Related work & Discussion
The evolution of the fixed-points of the exact score was studied in (Raya and Ambrogioni,, 2024) for the analysis of the spontaneous symmetry phenomenon breaking and in (Biroli and Mézard,, 2023) for the analysis of memorization and glassy phase transitions. The use of spectral gaps to quantify the dimensionality of the manifold was introduced in (Stanczuk et al.,, 2022), where the total manifold gap is analyzed. Several recent studies investigated the local linear structure of trained diffusion models. For example, (Kadkhodaie et al.,, 2024) studied the expansion of the Jacobian of trained models and described it as an optimal geometry-adaptive Harmonic representation. Similarly, (Chen et al., 2024b, ) characterized the linear expansion of the Jacobian of trained networks and characterized the resulting components in terms of their frequency content. Our work can be seen as a theoretical complement to this more applied line of research, as we provide a comprehensive random-matrix analysis of the phenomenon in tractable models. The dynamic geometry of diffusion manifolds was also investigated in (Chen et al., 2024a, ) using techniques inspired by research on latent generators such as GANs and autoencoders. Another recent work (Wang et al.,, 2024) uses parameterized low-rank score functions to show an equivalence between training diffusion model and subspace clustering. In our work, we prove results for the exact score for high-dimensional datasets. Finally, (Sakamoto et al.,, 2024) studies the dynamic geometry of tubular neighborhoods of the latent manifold and connect their singularities with spontaneous symmetry breaking events. Generative diffusion models exhibit rich geometric structures that have the potential to explain their impressive generative capabilities. Our work introduces the use of random-matrix theory techniques for the analysis of their dynamic local geometry and paves the way for the use of advanced statistical physics techniques, which may in the future unveil more global, topological and non-linear aspects of the dynamic geometry of diffusion generative models.

### v6

@@ 1 Introduction
Generative diffusion models have revolutionized the fields of computer vision and generative modeling, achieving state-of-the-art performance on image generation (Ho et al.,, 2020; Song and Ermon,, 2019; Yang et al.,, 2021) and video generation (Ho et al.,, 2022; Singer et al.,, 2022; Blattmann et al.,, 2023; Tim et al.,, 2024). Generative diffusion models synthesize images through a stochastic dynamical denoising process. Experimental and theoretical arguments suggest that different features such as frequency modes and class labels are generated at different times during the process. For example, it has been shown that separation between isolated classes, as in the case of mixture of Gaussian models, happens at critical phase transition points of spontaneous symmetry breaking (speciation events) (Biroli et al.,, 2024). It is also well known that subspaces corresponding to different frequency modes emerge at different times of diffusion (Kingma and Gao,, 2024). This idea has been recently refined by (Kadkhodaie et al.,, 2024), who showed that diffusion models give rise to a local decomposition of the image manifold into a basis of geometry-adaptive harmonic basis functions. These decomposition phenomena cannot be directly explained in terms of critical phase transitions as they are fundamentally linear processes. In this paper, we will provide a precise theoretical analysis of the separation of subspaces for data defined on low dimensional linear manifolds.
@@ 3 Background on generative diffusion models
where $\text{d}\boldsymbol{Z}_{t}$ is a standard Brownian motion. The formal solution of Eq. (3) can be given in terms of the heat kernel $p_{t}(\boldsymbol{x}_{t})=\mathbb{E}_{\boldsymbol{x}_{0}\sim p_{0}}\!\left[\frac{1}{\sqrt{2\pi t}}e^{-\frac{\left\lVert\boldsymbol{x}_{t}-\boldsymbol{x}_{0}\right\rVert_{2}^{2}}{2t}}\right].$ The target distribution $p_{0}(\boldsymbol{x})$ is then recovered by reversing the diffusion process (Anderson,, 1982). We initialize this reverse process from $\boldsymbol{x}_{t_{f}}\sim\mathcal{N}\left(0,t_{f}I_{d}\right)$ at some large time $t_{f}$, which evolves backward in time according to
@@ 5.4 The geometric phases and manifold overfitting
The probability density of data defined on a manifold is a spiked object $\delta_{\mathcal{M}}(\boldsymbol{x})\rho(\boldsymbol{x})$, where the Dirac-delta $\delta_{\mathcal{M}}(\boldsymbol{x})$ determines the manifold and $\rho(\boldsymbol{x})$ determines its internal density. Likelihood-based generative models are defined by a highly parameterized likelihood function $f(\boldsymbol{x};\boldsymbol{\theta})$, whose parameters are trained by minimizing the loss
@@
which maximizes the probability of the data given the model. This maximum likelihood loss is minimized if $f(\boldsymbol{x};\boldsymbol{\theta})=p_{0}(\boldsymbol{x})$. A trained likelihood-based model can only fit the true density by having it to diverge to infinity on the manifold. Such a divergence makes it impossible to correctly model the internal density $\rho(\boldsymbol{x})$. More problematically, the optimization problem becomes almost insensitive to the internal density $\rho(\boldsymbol{x})$. This phenomenon is called manifold overfitting (Loaiza-Ganem et al.,, 2022), since the trained model fits the manifold while ignoring its internal density, resulting in poor generation.
Our analysis suggests that the temporal dynamics of generative diffusion models overcome this limitation because, for intermediate values of $t$, the score is still sensitive to the density internal to the manifold, which can be identified through the differences in the tangent singular values. During this manifold coverage phase, the score directs the dispersion of the particles according to these differences, with higher singular values resulting in larger ‘opposing force’ from the score, which results in smaller displacements of the generated samples along these directions. For $t$ tending to zero, these differences are suppressed due to the divergence of the likelihood, which results in a score function that is orthogonal to the manifold and that is insensitive to $\rho(\boldsymbol{x})$. However, at this stage of generative diffusion the internal dispersion of the particles have already been affected by the previous coverage phase and therefore the manifold overfitting of the score does not negatively affect generation. Instead, the consolidation phase plays the important role of projecting the particles to the support of the data.
@@ 6.1 Linear manifolds
In the following, we provide an outline of our theoretical results on the distribution of spectral gaps in the matrix $J_{t}$ under random linear manifolds. This choice reflects the fact that the distribution and support of the data are usually not known in advance, and it is therefore important to quantify the statistical variability induced by this uncertainty. We will consider two different distributions for the random projection matrices $F$: an isotropic Gaussian case, and a multiple-variance one. To ensure tractability, we perform the analysis in the limit of large $d$ (visible) and $m$ (latent) dimensions while keeping the ratio $\alpha_{m}=m/d$ constant.
@@ 6.3 Intermediate gaps and subspaces with different variances
This same analysis can be extended to the more general case where the spectral density is known to be formed by different detached bulks, associated with hierarchically smaller variances of the data. The evolution of the intermediate gaps in a double-variance diffusion model is reported in Fig. 2: notice that $t_{\text{max}}=\mathcal{O}(\sigma_{2})$ is consistent with Fig. 2b and 2f, where the gap was found to be maximum in width. It is worth noting that subspaces with higher variances are the first ones to be explored by diffusion, and to be learned by the model. This point suggests that the model is sensitive to the parameters of the probability distribution on the manifold, as recently suggested by other works in the literature (see Section 9 for further details).
@@ 8 Experiments with natural image datasets
Trivial phase: at large times (i.e. from $t=200$ to $100$) the ordered spectrum of the singular values appears flat, suggesting the diffusive motion to be Brownian in the ambient space.
@@ 9 Related work & Discussion
The evolution of the fixed-points of the exact score was studied in (Raya and Ambrogioni,, 2024) for the analysis of the spontaneous symmetry phenomenon breaking and in (Biroli and Mézard,, 2023) for the analysis of memorization and glassy phase transitions. The use of spectral gaps to quantify the dimensionality of the manifold was introduced in (Stanczuk et al.,, 2022), where the total manifold gap is analyzed. Several recent studies investigated the local linear structure of trained diffusion models. For example, (Kadkhodaie et al.,, 2024) studied the expansion of the Jacobian of trained models and described it as an optimal geometry-adaptive Harmonic representation. Similarly, (Chen et al., 2024b, ) characterized the linear expansion of the Jacobian of trained networks and characterized the resulting components in terms of their frequency content. Our work can be seen as a theoretical complement to this more applied line of research, as we provide a comprehensive random-matrix analysis of the phenomenon in tractable models. The dynamic geometry of diffusion manifolds was also investigated in (Chen et al., 2024a, ) using techniques inspired by research on latent generators such as GANs and autoencoders. Another recent work (Peng et al.,, 2024) uses parameterized low-rank score functions to show an equivalence between training diffusion model and subspace clustering. In our work, we prove results for the exact score for high-dimensional datasets. Finally, (Sakamoto et al.,, 2024) studies the dynamic geometry of tubular neighborhoods of the latent manifold and connect their singularities with spontaneous symmetry breaking events. Generative diffusion models exhibit rich geometric structures that have the potential to explain their impressive generative capabilities. Our work introduces the use of random-matrix theory techniques for the analysis of their dynamic local geometry and paves the way for the use of advanced statistical physics techniques, which may in the future unveil more global, topological and non-linear aspects of the dynamic geometry of diffusion generative models.
Acknowledgements
This publication benefited from European Union - Next Generation EU funds, component M4.C2, investment 1.1. - CUP J53D23001330001.